How do you do y= 2x + 3 using substitution?
y= 3x + 1

Answer by checkley71(8403)

y=2x+3 now we substitute 2x+3 for y in equation #2 we get
2x+3=3x+1 or -x=-2 or x=2 then y=2*2+3 or y=7 & y=2*3+1 or y=7